# ContraGen: A Multi-Agent Generation Framework for Enterprise Contradictions Detection

### v1

@@ A-A Prompts for Data Generation
A-B Prompts for Contradiction Injection
@@
Prompt: Generate Contradiction You are helping generate business documents that contain realistic internal contradictions. You will be given the original document and one important sentence selected from it (the target statement). Your task: Write 1 or 2 sentences that DIRECTLY CONTRADICTS the target statement after selecting one of the contradiction types mentioned in the few-shot examples. The contradiction should be HARD and DIRECT, not soft or hedging. Use the same tone and style as the original document. Do NOT use words like ‘however’, ‘while’, ‘although’, ‘but’, ‘may’, ‘might’, ‘could’, ‘sometimes’, ‘certain’, ‘extended’, ‘flexibility’. Make a clear, confident business statement that contradicts the target. The contradiction should be realistic, specific, and plausible within a formal enterprise setting. The paragraph should sound like a natural part of the document, not an exception or explanation. FEW-SHOT EXAMPLES: TARGET: The new performance evaluation cycle begins on January 15, 2024.
CONTRADICTION: The official launch of the new evaluation cycle will be in late Q1 accommodating inter-departmental calibration sessions.
TYPE: temporal
TARGET: The North America team exceeded Q2 revenue targets by \$12M.
CONTRADICTION: The North America team fell short of Q2 revenue by \$5M.
TYPE: numerical
TARGET: This policy is issued under the authority of the Global Compliance Office.
CONTRADICTION: This policy is issued by the Strategy Operations Unit.
TYPE: authority
TARGET: All employee leave requests must be submitted via the HR portal.
CONTRADICTION: All employee leave requests must be submitted through departmental admins.
TYPE: process
TARGET: Remote work is mandatory for Tier 1 engineers.
CONTRADICTION: Remote work is not permitted for any Tier 1 engineering staff.
TYPE: policy reversal
TARGET: The new training policy applies to all regional offices globally.
CONTRADICTION: The new training policy applies to the APAC offices.
TYPE: specificity
TARGET STATEMENT: {target_statement} ORIGINAL DOCUMENT: {document_text}
@@
A-C Prompt used for LLM Contradiction Judge

### v2

@@ A.1 Prompts for Data Generation
A.2 Prompts for Contradiction Injection
@@
Prompt: Generate Contradiction You are helping generate business documents that contain realistic internal contradictions. You will be given the original document and one important sentence selected from it (the target statement). Your task: Write 1 or 2 sentences that DIRECTLY CONTRADICTS the target statement after selecting one of the contradiction types mentioned in the few-shot examples. The contradiction should be HARD and DIRECT, not soft or hedging. Use the same tone and style as the original document. Do NOT use words like ‘however’, ‘while’, ‘although’, ‘but’, ‘may’, ‘might’, ‘could’, ‘sometimes’, ‘certain’, ‘extended’, ‘flexibility’. Make a clear, confident business statement that contradicts the target. The contradiction should be realistic, specific, and plausible within a formal enterprise setting. The paragraph should sound like a natural part of the document, not an exception or explanation. FEW-SHOT EXAMPLES: {few_shot_examples} TARGET STATEMENT: {target_statement} ORIGINAL DOCUMENT: {document_text}
@@
A.3 Prompt used for LLM Contradiction Judge